OpenStudy (anonymous):

8.5 gallons of gas to drive 243.1 miles, 3 gallons of gas to drive __ miles.

OpenStudy (mathstudent55):

This is a proportion problem. Solve for x. Cross multiply. \(\dfrac{8.5}{243.1} = \dfrac{3}{x} \)

OpenStudy (anonymous):

i thought you had to find the slope \[m =\frac{ {y2-y1} }{ x2-x1 }\]